OpenStudy (anonymous):

what would the slope be for the equation 4x - 3y = 12?

OpenStudy (anonymous):

first put it in the form of y=mx+b -3y=-4x+12 y=4/3x -4 the slope is m which is 4/3

OpenStudy (anonymous):

so move to the side 4 and up 3?

OpenStudy (anonymous):

move up4 and over 3